OpenStudy (anonymous):

A 15 kw motor provides power for an elevator of a building. What is the minimum time needed for the elevator to rise 30m when its total mass when loaded in 900kg?

OpenStudy (anonymous):

\[t = \frac{ w }{ P }=\frac{ mgh }{ P }\]

OpenStudy (anonymous):

@ganeshie8 @hartnn @Abhisar @Nurali I got t = 17,640sec

OpenStudy (anonymous):

\[t = \frac{ 900kg * 9.8 m/\sec ^{2}* 30m}{ 15kw }\]

OpenStudy (anonymous):

Is your answer correct?

OpenStudy (anonymous):

I'm asking you all. Is it correct?

OpenStudy (anonymous):

I got 17.66seconds. You shall be correct then.

OpenStudy (anonymous):

Fisry I did P/F=v then d/v=t

OpenStudy (anonymous):

Can you please show your work just so I can understand it more...?

OpenStudy (anonymous):

Of course.

OpenStudy (anonymous):

P/F=v or 15000/900*9.81= 1.698m/s t=d/v=30/1.698=17.667s.

OpenStudy (anonymous):

One second. I'm trying to understand this... Does it still use the same formula I provided above? I have to use it

OpenStudy (anonymous):

Yes of course. Your formula is also correct and is much more direct actually or let me say, more accurate, because you made t subject of the formula and use given values in one single method. I combined Power formula with that of velocity from motion theory.